OpenStudy (anonymous):

if m varies directly as p, and m=5 when p=7, what is the constant of variation?

OpenStudy (anonymous):

m varies directly as p means m = kp

OpenStudy (anonymous):

Substituting in for m and p 5 = k(7)

OpenStudy (anonymous):

Solve for k by dividing by 7 5 = k(7) - --- 7 7 5/7 = k

OpenStudy (anonymous):

Now start over with m = kp but put in 5/7 for k m = 5/7p

OpenStudy (anonymous):

Oops.. didn't read the problem.. your constant of variation means the same as the value of k so the answer is 5/7

OpenStudy (anonymous):

haha thank you :)